Topstone Industries has an expected EBIT of $1,000,000, a cost of equity of 11% and a cost of debt of 5%. Topstone s debt-to-equity ratio is 1/3. The corporate tax rate is 25%. What is the appropriate discount rate to be used under the APV method to value Topstone?
Answer #1

It can be measured through weighted cost of capital so-

Weight of debt= 1/(1+3)= 25%

Weight of equity= 3/(3+1)=75%

Cost of debt before tax= 5%

Cost of equity= 11%

Tax rate = 25%

WACC= (cost of equity × weight of equity)+ (cost of debt×weight of debt)(1-tax rate)

= (13×.75)+(5×.25)(1-.25)

= (9.75+.9375)

= 10.6875%

So appropriate discount rate to be used= 10.6875%